factor the polynomial completely.

2b^4+14b^3-16b-112

Answer by Alan3354(69443)

factor the polynomial completely.
2b^4+14b^3-16b-112
= 2*(b^4 + 7b^3 - 8b - 56)
= 2*(b+7)*(b^3 - 8)
= 2*(b+7)*(b-2)*(b^2 + 2b + 4)